### v1

@@ Abstract
The use of model order reduction techniques in combination with ensemble-based methods for estimating the state of systems described by nonlinear partial differential equations has been of great interest in recent years in the data assimilation community. Methods such as the multi-fidelity ensemble Kalman filter (MF-EnKF) and the multi-level ensemble Kalman filter (ML-EnKF) are recognized as state-of-the-art techniques. However, in many cases, the construction of low-fidelity models in an offline stage, before solving the data assimilation problem, prevents them from being both accurate and computationally efficient. In our work, we investigate the use of adaptive reduced basis techniques in which the approximation space is modified online based on the information that is extracted from a limited number of full order solutions and that is carried by the past models. This allows to simultaneously ensure good accuracy and low cost for the employed models and thus improve the performance of the multi-fidelity and multi-level methods.
@@ 1 Introduction
In our study, we restrict our focus to parabolic dissipative problems with sufficient regularity. Furthermore, we consider well-behaved initial uncertainty distributions which can be effectively approximated with low-dimensional spaces. Within these distributions, we encode the problem’s regularity, or, when available, some information on the long time behavior of the system.
The paper is organized as follows. In Section 2, we briefly introduce the filtering problem alongside the classical Bayesian and Kalman filters. Then, in Section 3, we offer a brief overview of the technique for reduced order modeling, focusing specifically on proper orthogonal decomposition (POD) and Galerkin projections. Section 4 is dedicated to a detailed discussion of the ML-EnKF and MF-EnKF, with Section 4.3 specifically focused on the innovative elements introduced with our approach. Finally, in Section 5, we test the proposed approaches on the quasi-geostrophic equations (QGE) and we compare their performances for different surrogate models. In Section 6 we draw conclusions on the proposed method and on its numerical performance.
@@ 2 The filtering problem
The ability to propagate this uncertainty from one point in time to another is also crucial for addressing the filtering problem. In fact, along with the assimilation of new data, it is essential to evolve the state uncertainty consistently with the model expected to describe the evolution of the ‘true’ state. In our case, we assume that the random state, $\omega(t_{k})\in\mathcal{V}$, evolves from $t_{k}$ to $t>t_{k}$ via the flow $\Psi:\mathcal{V}\times\mathbb{R}\rightarrow\mathcal{V}$, so that $\omega(t)=\Psi(\omega(t_{k}),t-t_{k})$. The forecast of the state over each assimilation window $\mathcal{I}_{k}$ is therefore obtained as
$\omega(t_{k+1})=\Psi(\omega(t_{k}),t_{k+1}-t_{k}),\qquad\forall k\in\left[0,\ldots,N_{T}\right].$ (1)
For systems governed by partial differential equations, the forecast model $\Psi(\,\cdot\,,t_{k+1}-t_{k})$ is obtained by integrating a time-dependent PDE over the assimilation window $\mathcal{I}_{k}$. The general parabolic partial differential equation considered in this work can be formulated in weak form as follows: given the initial condition $\omega(t_{k})\in\mathcal{V}$, determine $\omega(t_{k+1})\in\mathcal{V}$ such that
@@
To describe the Bayesian filter we extend the transformation $\Psi(\,\cdot\,,t_{k+1}-t_{k})$ to the set of the probability distributions on the measurable space $(\mathcal{V},\mathcal{B}(\mathcal{V}))$ by introducing the pushforward operator $\Psi_{\#}(\,\cdot\,,t_{k+1}-t_{k})$. The Bayesian filter sarkka2013 , which represents a general probabilistic approach to the filtering problem, can be written in terms of this operation and of Bayesian updates, as in Algorithm 1.
@@
(i) Analysis step: update the probability distribution according to Bayes’ rule
@@
(ii) Prediction step: predict the distribution at the next assimilation time
$\pi(\omega(t_{k+1})|\mathbf{d}_{k},\mathbf{d}_{k-1},...)=\Psi_{\#}(\pi(\omega(t_{k})|\mathbf{d}_{k},\mathbf{d}_{k-1},...),t_{k+1}-t_{k}).$
In practice, this filter is impossible to implement in most cases and it must be substituted by approximate filters that may be more practical but possibly less effective. In this work, we focus on the EnKF and on its variants. This family of filters relies on two fundamental assumptions. First, we assume that the probability distributions characterizing the uncertainty on $\omega^{\star}(t)$ can be approximated by an empirical distribution involving a discrete set of states. Second, we further assume that in the analysis step, all the distributions can be treated as Gaussians and thus, the Kalman update formula can be applied. The steps of the EnKF and its variants are summarized in Algorithm 2 and then detailed in Section 2.1 and 2.2.
@@
  1. (0)
    Initialization: Generate the initial ensembleby sampling states from $\pi(\omega(t_{0}))$.
@@ 2.1 Choosing the initial ensembles
$E_{k|r}^{\Diamond}\coloneqq\{\omega_{k|r}^{\Diamond,n}\in\mathcal{V}\;|\;\omega_{k|r}^{\Diamond,n}\sim\pi(\omega(t_{k})|\mathbf{d}_{r})\}_{n=1}^{N_{\Diamond}},$
where $0\leq k\leq N_{T}$ and $-1\leq r\leq k$. The sampling probability distribution $\pi(\omega(t_{k})|\mathbf{d}_{r})$ corresponds to the uncertainty measure that we aim to approximate using the ensemble. Throughout, we adopt the following convention for the ensemble members: the superscript $\Diamond$ denotes the ensemble family, and the index within the ensemble family, i.e., “$\Diamond,n$” refers to the $n$th state in the $\Diamond$ ensemble; the subscripts “$k|r$” refer to the approximation time $t_{k}$ and the data assimilation time $t_{r}$, i.e., $k|r$ refers to an approximation of $\omega^{\star}{(t_{k})}$ using data updated at time $t_{r}$. The initial ensemble, which we denote by $E_{0|-1}^{\Diamond}$, reflects a partial knowledge of the ‘true’ state $\omega^{\star}(t_{0})$ before any experimental data is made available. In this work, we consider two alternative sampling distributions for the ensemble $E_{0|-1}^{\Diamond}$, both arising from the knowledge of the model describing the dynamical system. The first option is based on the sole knowledge of the regularity of the state, e.g., in smooth dissipative processes; the second option assumes that the dynamics is fully characterized. For example, if the dynamical system possesses an attractor and the ‘true’ state lies within a basin of attraction, the associated invariant measure can be employed as a sampling distribution. In the numerical experiments of Section 5, we show how these two possible choices affect the reconstruction of the state.
@@ 2.2 The ensemble Kalman filter
The EnKF algorithm starts with the initialization of an initial ensemble, $E^{P}_{0|-1}$, of size $N_{P}\in\mathbb{N}^{+}$, whose states are independently sampled according to the rules described in Section 2.1.
At the $k$th iteration, with $k\in\{0,\ldots,N_{T}\}$, the first step of the EnKF is the analysis or update step (Algorithm 2, Step (i)) performed when new experimental data are made available. This assimilation step relies on the Kalman update formula to integrate the noisy data $\mathbf{d}_{k}$ with synthetic observations of the predicted states in the forecast ensemble $E^{P}_{k|k-1}$. By assuming an ensemble of states in $\mathcal{V}$, the update step can be stated as follows: for all $n\in\{1,...,N_{P}\}$,
@@
where $\mathbf{K}_{k}^{\text{E}}:\mathcal{Y}\rightarrow\mathcal{V}$ is the Kalman gain evensen2003 defined as $\mathbf{K}_{k}^{\text{E}}\coloneqq\mathbf{Q}_{k}^{\text{E}}(\mathbf{P}_{k}^{\text{E}}+\bm{\Sigma})^{-1}$. The correlation operators $\mathbf{Q}_{k}^{\text{E}}:\mathcal{Y}\rightarrow\mathcal{V}$ and $\mathbf{P}_{k}^{\text{E}}:\mathcal{Y}\rightarrow\mathcal{Y}$ are the best unbiased estimators for the state-measurement covariance and for the measurement-measurement covariance, respectively, and can be computed as
@@
The analysis is followed by the prediction step (Algorithm 2, Step (ii)) in which all the states in the post-analysis ensemble $E^{P}_{k|k}$ are advanced over the assimilation window $\mathcal{I}_{k}$ using the high-fidelity model $\Psi(\,\cdot\,,t_{k+1}-t_{k})$, namely
$\omega^{P,n}_{k+1|k}=\Psi(\omega^{P,n}_{k|k},t_{k+1}-t_{k}),\qquad\forall\,n\in\{1,...,N_{P}\}.$ (3)
Solving equation (3) for all ensemble members has arithmetic complexity $\mathcal{O}(\alpha(\mathcal{N})N_{P})$, where $\alpha(\mathcal{N})$ denotes the cost of advancing a single state one assimilation step in the future. This is usually performed by solving the PDE (2) $N_{P}$ times and can become prohibitively expensive whenever $\mathcal{N}$ is large. Hence, this step can become the primary cost in the filtering process due to its dependence on $\mathcal{N}$ and the requirement of a sufficiently large ensemble size $N_{P}$ to achieve a good data assimilation accuracy. In this context, the use of surrogate models and reduced order modeling thus becomes crucial.
@@ 3 Reduced order modeling
$\Xi\coloneqq\{\Psi(\omega(t_{k}),t-t_{k}),\;\forall t\in\mathcal{I},\,\omega(t_{k})\sim\pi(\omega(t_{k})|\mathbf{d}_{r})\},$ (4)
with $-1\leq r\leq k-1$, obtained by solving the partial differential equation (2) over the interval $\mathcal{I}=[t_{k},t_{k}+T]$ of length $T\in\mathbb{R}^{+}$. The problem is considered reducible if the solution set can be accurately approximated by a low-dimensional linear subspace of $\mathcal{V}$, i.e., if, for a given tolerance $\varepsilon\in\mathbb{R}^{+}$, there exists a subspace $\mathcal{V}_{\varepsilon}\subset\mathcal{V}$ of size $N_{\varepsilon}\ll\mathcal{N}$ such that
$J(\mathcal{V}_{\varepsilon})\coloneqq\dfrac{1}{T}{\int_{\mathcal{I}}\int_{\mathcal{V}}\|(\text{Id}-\Pi_{\mathcal{V}_{\varepsilon}})\Psi(\omega(t_{k}),t-t_{k})\||^{2}_{\mathcal{V}}\,\pi(d\omega(t_{k})\mathbf{d}_{r})\,dt}\leq\varepsilon,$ (5)
where $\Pi_{\mathcal{V}_{\varepsilon}}:\mathcal{V}\rightarrow\mathcal{V}_{\varepsilon}$ denotes the $\mathcal{V}$-orthogonal projection onto $\mathcal{V}_{\varepsilon}$ and $\text{Id}:\mathcal{V}\rightarrow\mathcal{V}$ denotes the identity operator. Since one typically does not have complete information about $\omega(t_{k})$ and its probability distribution, an empirical approximation of the integral above is needed to select a suitable subspace $\mathcal{V}_{\varepsilon}$, as it will be discussed in Section 3.1.
@@
where the operator $\mathcal{F}_{\varepsilon}$ is obtained by projecting $\mathcal{F}$ from (2) onto the reduced space $\mathcal{V}_{\varepsilon}$. If we consider the problem over the temporal interval $\mathcal{I}_{k}=[t_{k},t_{k+1}]$, we can introduce the reduced flux $\Psi_{\varepsilon}(\,\cdot\,,t_{k+1}-t_{k}):\mathcal{V}_{\varepsilon}\rightarrow\mathcal{V}_{\varepsilon}$ so that
$\omega_{\varepsilon}(t_{k+1})=\Psi_{\varepsilon}(\omega_{\varepsilon}(t_{k}),t_{k+1}-t_{k}).$
Assuming that the full order problem is reducible over the interval $\mathcal{I}_{k}$ given the prior $\pi(\omega(t_{k})|\mathbf{d}_{k})$, the reduced model (6) can be solved at a cost independent of the size $\mathcal{N}$ of the full order problem under the assumption of linearity and parameter-separability of the operator $\mathcal{F}$. Addressing general non-linear operators may necessitate the application of hyper-reduction techniques, such as empirical interpolation (EIM) barrault2004 , linear program empirical quadrature yano2019 , or empirical cubature hernandez2017 .
@@
We propose an adaptive strategy which entails that the training phase is no longer confined to an offline phase, as in classical reduced basis methods. Instead, we employ online generated snapshots coming from the data assimilation algorithm itself. This approach substantially reduces the computational costs but also introduces extra challenges, which will be discussed and addressed in Section 4.3.

### v2

@@ Abstract
The use of reduced order modeling techniques in combination with ensemble-based methods for estimating the state of systems described by nonlinear partial differential equations has been of great interest in recent years in the data assimilation community. Methods such as the multi-fidelity ensemble Kalman filter (MF-EnKF) and the multi-level ensemble Kalman filter (ML-EnKF) are recognized as state-of-the-art techniques. However, in many cases, the construction of low-fidelity models in an offline stage, before solving the data assimilation problem, prevents them from being both accurate and computationally efficient. In our work, we investigate the use of adaptive reduced basis techniques in which the approximation space is modified online by combining information extracted from a limited number of full order solutions and information extracted from reduced models trained at previous time steps. This allows to simultaneously ensure good accuracy and low cost for the employed models and thus improve the performance of the multi-fidelity and multi-level methods.
@@ 1 Introduction
In our study, we focus on parabolic dissipative problems with sufficiently regular solutions. These hypotheses offer an extensively studied framework for the use of reduced-order modeling techniques and ensure a good theoretical understanding of the convergence of data assimilation algorithms oljava2018 . Furthermore, we consider initial uncertainty distributions that can be effectively approximated within low-dimensional spaces with defined moments and sufficiently light tails, a necessary condition to ensure the problem’s reducibility. Within these distributions, we encode the problem’s regularity or, when available, some information on the system’s long-term behavior.
The paper is organized as follows. In Section 2, we briefly introduce the filtering problem alongside the classical Bayesian and Kalman filters. Then, in Section 3, we offer a brief overview of the technique for reduced order modeling, focusing specifically on proper orthogonal decomposition (POD) and Galerkin projections. Section 4 is dedicated to a detailed discussion of the ML-EnKF and MF-EnKF, with Subsection 4.3 specifically focused on the innovative elements introduced with our approach. Finally, in Section 5, we test the proposed approaches on the quasi-geostrophic equations (QGE) and we compare their performances for different surrogate models. In Section 6 we draw conclusions on the proposed method and on its numerical performance.
@@ 2 The filtering problem
The ability to propagate this uncertainty from one point in time to another is also crucial for addressing the filtering problem. In fact, along with the assimilation of new data, it is essential to evolve the state uncertainty consistently with the model expected to describe the evolution of the ‘true’ state. In our case, we assume that the random state, $\omega(t_{k})\in\mathcal{V}$, evolves from $t_{k}$ to $t>t_{k}$ via the flow $\Psi:\mathcal{V}\times\mathbb{R}\times\mathbb{R}\rightarrow\mathcal{V}$, so that $\omega(t)={\color[rgb]{0,0,0}\Psi(\omega(t_{k}),t_{k},t)}$. The forecast of the state over each assimilation window $\mathcal{I}_{k}$ is therefore obtained as
$\omega(t_{k+1})={\color[rgb]{0,0,0}\Psi(\omega(t_{k}),t_{k},t_{k+1})},\qquad\forall k\in\left[0,\ldots,N_{T}\right].$ (1)
For systems governed by partial differential equations, the forecast model $\Psi(\,\cdot\,,t_{k},t_{k+1})$ is obtained by integrating a time-dependent PDE over the assimilation window $\mathcal{I}_{k}$. The general parabolic partial differential equation considered in this work can be formulated in weak form as follows: given the initial condition $\omega(t_{k})\in\mathcal{V}$, determine $\omega(t_{k+1})\in\mathcal{V}$ such that
@@
To describe the Bayesian filter we extend the transformation ${\color[rgb]{0,0,0}\Psi(\,\cdot\,,t_{k},t_{k+1})}$ to the set of the probability distributions on the measurable space $(\mathcal{V},\mathcal{B}(\mathcal{V}))$ by introducing the pushforward operator ${\color[rgb]{0,0,0}\Psi_{\#}(\,\cdot\,,t_{k},t_{k+1})}$. The Bayesian filter sarkka2013 , which represents a general probabilistic approach to the filtering problem, can be written in terms of this operation and of Bayesian updates, as in Algorithm 1.
@@
  1. (i)
    Analysis step: update the probability distribution according to Bayes’ rule
@@
  2. (ii)
    Prediction step: predict the probability distribution at time $t_{k+1}$
    $\pi(\omega(t_{k+1})|\mathbf{d}_{k},\mathbf{d}_{k-1},...)={\color[rgb]{0,0,0}\Psi_{\#}(\pi(\omega(t_{k})|\mathbf{d}_{k},\mathbf{d}_{k-1},...),t_{k},t_{k+1})}.$
In practice, this filter is impossible to implement in most cases and it must be substituted by approximate filters that may be more practical but possibly less effective. In this work, we focus on the EnKF and on its variants. This family of filters relies on two fundamental assumptions. First, we assume that the probability distributions characterizing the uncertainty on $\omega^{\star}(t)$ can be approximated by an empirical distribution involving a discrete set of states. Second, we further assume that in the analysis step, all the distributions can be treated as Gaussians and thus, the Kalman update formula can be applied. The steps of the EnKF and its variants are summarized in Algorithm 2 and then detailed in Subsection 2.1 and 2.2.
@@
(0) Initialization: Generate the initial ensemble by sampling states from $\pi(\omega(t_{0}))$.
@@ 2.1 Choosing the initial ensembles
$E_{k|r}^{\Diamond}:=\{\omega_{k|r}^{\Diamond,n}\in\mathcal{V}\;|\;\omega_{k|r}^{\Diamond,n}\sim\pi(\omega(t_{k})|\mathbf{d}_{r})\}_{n=1}^{N_{\Diamond}},$
where $0\leq k\leq N_{T}$ and $-1\leq r\leq k$. The sampling probability distribution $\pi(\omega(t_{k})|\mathbf{d}_{r})$ corresponds to the uncertainty measure that we aim to approximate using the ensemble. Throughout, we adopt the following convention for the ensemble members: the superscript $\Diamond$ denotes the ensemble family, and the index within the ensemble family, i.e., "$\Diamond,n"$ refers to the $n$th state in the $\Diamond$ ensemble; the subscripts "$k|r"$ refer to the approximation time $t_{k}$ and the data assimilation time $t_{r}$, i.e., $k|r$ refers to an approximation of $\omega^{\star}{(t_{k})}$ using data updated at time $t_{r}$. The initial ensemble, which we denote by $E_{0|-1}^{\Diamond}$, reflects a partial knowledge of the ‘true’ state $\omega^{\star}(t_{0})$ before any experimental data is made available. In this work, we consider two alternative sampling distributions for the ensemble $E_{0|-1}^{\Diamond}$, both arising from the knowledge of the model describing the dynamical system. The first option is based on the sole knowledge of the regularity of the state, e.g., in smooth dissipative processes; the second option assumes that the dynamics is fully characterized. For example, if the dynamical system possesses an attractor and the ‘true’ state lies within a basin of attraction, the associated invariant measure can be employed as a sampling distribution. In practice, the most common strategy to obtain random states distributed according to this measure involves simulating the system’s dynamics for a sufficiently long time and then randomly selecting states from this long-time trajectory. In the numerical experiments of Section 5, we show how these two possible choices affect the reconstruction of the state.
@@ 2.2 The ensemble Kalman filter
The EnKF algorithm starts with the initialization of an initial ensemble, $E^{P}_{0|-1}$, of size $N_{P}\in\mathbb{N}^{+}$, whose states are independently sampled according to the rules described in Subsection 2.1.
At the $k$th iteration, with $k\in\{0,\ldots,N_{T}\}$, the first step of the EnKF is the analysis or update step (Algorithm 2, Step (i)) performed when new experimental data are made available. This step relies on the Kalman update formula to integrate the noisy data $\mathbf{d}_{k}$ with synthetic observations of the predicted states in the forecast ensemble $E^{P}_{k|k-1}$. By assuming an ensemble of states in $\mathcal{V}$, the update step can be stated as follows: for all $n\in\{1,...,N_{P}\}$,
@@
where $\mathbf{K}_{k}^{\text{E}}:\mathcal{Y}\rightarrow\mathcal{V}$ is the Kalman gain evensen2003 defined as $\mathbf{K}_{k}^{\text{E}}:=\mathbf{Q}_{k}^{\text{E}}(\mathbf{P}_{k}^{\text{E}}+\bm{\Sigma})^{-1}$. The correlation operators $\mathbf{Q}_{k}^{\text{E}}:\mathcal{Y}\rightarrow\mathcal{V}$ and $\mathbf{P}_{k}^{\text{E}}:\mathcal{Y}\rightarrow\mathcal{Y}$ are the best unbiased estimators for the state-measurement covariance and for the measurement-measurement covariance, respectively, and can be computed as
@@
The analysis is followed by the prediction step (Algorithm 2, Step (ii)) in which all the states in the post-analysis ensemble $E^{P}_{k|k}$ are advanced over the assimilation window $\mathcal{I}_{k}$ using the high-fidelity model ${\color[rgb]{0,0,0}\Psi(\,\cdot\,,t_{k},t_{k+1})}$, namely
$\omega^{P,n}_{k+1|k}={\color[rgb]{0,0,0}\Psi(\omega^{P,n}_{k|k},t_{k},t_{k+1})},\qquad\forall\,n\in\{1,...,N_{P}\}.$ (3)
Solving equation (3) for all ensemble members has arithmetic complexity $\mathcal{O}(\alpha(\mathcal{N})N_{P})$, where $\alpha(\mathcal{N})$ denotes the cost of advancing a single state from time $t_{k}$ to time $t_{k+1}$. This is usually performed by solving the PDE (2) $N_{P}$ times and can become prohibitively expensive whenever $\mathcal{N}$ is large. Hence, this step can become the primary cost in the filtering process due to its dependence on $\mathcal{N}$ and the requirement of a sufficiently large ensemble size $N_{P}$ to achieve a good data assimilation accuracy. In this context, the use of surrogate models and reduced order modeling thus becomes crucial.
@@ 3 Reduced order modeling
$\Xi:=\{{\color[rgb]{0,0,0}\Psi(\omega(t_{k}),t_{k},t)},\;\forall t\in\mathcal{I},\,\omega(t_{k})\sim\pi(\omega(t_{k})|\mathbf{d}_{r})\},$ (4)
with $-1\leq r\leq k-1$, obtained by solving the PDE (2) over the interval $\mathcal{I}=[t_{k},t_{k}+T]$ of length $T\in\mathbb{R}^{+}$. The problem is considered reducible if the solution set can be accurately approximated by a low-dimensional linear subspace of $\mathcal{V}$, i.e., if, for a given tolerance $\varepsilon\in\mathbb{R}^{+}$, there exists a subspace $\mathcal{V}_{\varepsilon}\subset\mathcal{V}$ of size $N_{\varepsilon}\ll\mathcal{N}$ such that
[…]
where $\Pi_{\mathcal{V}_{\varepsilon}}:\mathcal{V}\rightarrow\mathcal{V}_{\varepsilon}$ denotes the $\mathcal{V}$-orthogonal projection onto $\mathcal{V}_{\varepsilon}$ and $\text{Id}:\mathcal{V}\rightarrow\mathcal{V}$ denotes the identity operator. Since one typically does not have complete information about $\omega(t_{k})$ and its probability distribution, an empirical approximation of the integral above is needed to select a suitable subspace $\mathcal{V}_{\varepsilon}$, as it will be discussed in Subsection 3.1.
@@
where the operator $\mathcal{F}_{\varepsilon}$ is obtained by projecting $\mathcal{F}$ from (2) onto the reduced space $\mathcal{V}_{\varepsilon}$. If we consider the problem over the temporal interval $\mathcal{I}_{k}=[t_{k},t_{k+1}]$, we can introduce the reduced flux ${\color[rgb]{0,0,0}\Psi_{\varepsilon}(\,\cdot\,,t_{k},t_{k+1})}:\mathcal{V}_{\varepsilon}\rightarrow\mathcal{V}_{\varepsilon}$ so that
$\omega_{\varepsilon}(t_{k+1})={\color[rgb]{0,0,0}\Psi_{\varepsilon}(\omega_{\varepsilon}(t_{k}),t_{k},t_{k+1})}.$
Assuming that the full order problem is reducible over the interval $\mathcal{I}_{k}$ given the prior $\pi(\omega(t_{k})|\mathbf{d}_{k})$, the reduced model (6) can be solved at a cost independent of the size $\mathcal{N}$ of the full order problem under the assumption of linearity and parameter-separability of the operator $\mathcal{F}$. Addressing general non-linear operators may necessitate the application of hyper-reduction techniques, such as empirical interpolation (EIM) barrault2004 , linear program empirical quadrature yano2019 , $\ell^{p}$-quasi-norm minimization manucci2022 , or empirical cubature hernandez2017 .
@@
The idea of using online adaptive reduced-order modeling techniques to lower the offline training costs and address problems with limited global reducibility is relevant and of current interest. Notable examples can be found in the fields of optimization carlberg2008 ; peherstorfer2015 ; qian2017 , simulation washabaugh2012 ; pagliantini2021 , and control kramer2017 of dynamical systems, as well as in parameter estimation kartmann2023 and uncertainty quantification vidlickova2022 . Unlike existing methods, we neither use non-intrusive approaches nor develop evolution equations for the approximation spaces. Instead, we propose a discontinuous online adaptation of the reduced basis, in line with the filter’s discontinuous nature. Also, unlike trust region methods, we do not rely on error bounds for space inflation but use a POD-based approach. Additionally, we introduce a deflation phase to remove unused bases, thus adapting the size of the approximation space to the problem’s local reducibility.
We propose an adaptive strategy which entails that the training phase is no longer confined to an offline phase, as in classical reduced basis methods. Instead, we employ online generated snapshots coming from the data assimilation algorithm itself. This approach substantially reduces the computational costs but also introduces extra challenges, which will be discussed and addressed in Subsection 4.3.